OpenStudy (darkbluechocobo):

The formula for continuously compounded interest on a principal investment P at a given interest rate r over time t in years is given by A = Pe^rt. You have deposited $12,000 in an account that pays 4.15% interest, compounded continuously. How long will it take for your initial investment to grow six times its original worth? Round your answer to the nearest tenth.

OpenStudy (darkbluechocobo):

@satellite73 could you help me Im not so good with compound interest S:

OpenStudy (darkbluechocobo):

would this be 43.2 years?

OpenStudy (kropot72):

\[\large A=Pe^{rt}\ ..............(1)\] If we divide both sides of (1) by P we get: \[\large \frac{A}{P}=e^{rt}\ ............(2)\] In the question it is given that A/P = 6, and r expressed as a decimal is 0.0415. Plugging these values into (2) gives: \[\large 6=e^{0.0415t}\ ...........(3)\] Now you just need to solve (3) to find the value of t.

OpenStudy (darkbluechocobo):

See this is where it comes difficult for me Because I dont understand how to get an exponent by itself

OpenStudy (kropot72):

Take natural logs of both sides of (3).

OpenStudy (darkbluechocobo):

so if natural then would it be 6e?

OpenStudy (darkbluechocobo):

im kinda new to log S:

OpenStudy (darkbluechocobo):

nvm so ln6=0.0415t?

OpenStudy (kropot72):

Your equation: \[\large \ln 6=0.0415t\] is correct. Now you need to divide both sides of the equation by 0.0415 to find the value of t in years.